solve for n
n=575/8*2/9

Answer by Theo(13342)

no parentheses, so you go from left to right in the following order.
exponents first
multiplications and divides next
additions and subtractions next.

your equation is:

n = 575/8*2/9

all multiplications and divides so just do them from left to right.

575/8 = 71.875

71.875 * 2 = 143.75

143.75 / 9 = 15.9722222........

with parentheses to clearly identify the order, this would look like this:

n = 575/8*2/9 becomes ((575/8)*2)/9

you would do the inner parentheses first which would be 575/8.

you would do the next inner parentheses next which would be (575/8)*2

you would do the next inner parentheses next which would be ((575/8)*2)/9

in fraction form, the operation would look like this:

n = 575/8*2/9

575/8 = 575/8

575/8 * 2 = (575*2)/8

((575*2)/8)/9 = (575*2)/(8*9)

simplify this to get 575/(4*9) = 575/36

convert this to decimal form and you get 15.9722222......; same as above.